OpenStudy (anonymous):

2x+y=5 3x-3y=3 solve by substituting. it would be nnice to get some helpful tips please! thank you. i just need a method that could stick to me to find them.. it's mega confusing how my teacher explains it. & i know it's not that hard to learn.

OpenStudy (whpalmer4):

\[2x+y=5\]\[3x-3y=3\] Okay, pick one of the equations that is easy to arrange so one of the variables is in terms of the other. In this case, it is easy to rearrange the first equation to give y in terms of x.\[2x+y=5\]\[y=5-2x\] Now, everywhere in the other equation where you see y, instead write 5-2x: \[3x-3(5-2x)=3\]\[3x-15+6x=3\] (remember that -3 * (-2x) = 6x) \[9x=18\]\[x=2\] Any question about what I did so far?

OpenStudy (anonymous):

y=5-2x substitute into the second eqn an solve by yourself, i suggest

OpenStudy (whpalmer4):

Now go back to the first equation and plug in the value of x we found and solve for y. \[y=5-2(2) = 1\] Check your work by putting both x and y values into the equations and making sure they still work. 2(2)+1=5, 3(2)-3(1)=3

OpenStudy (anonymous):

x=2 y=1

OpenStudy (anonymous):

@whpalmer4: thank you so much for the help. i truly appreciate it.

OpenStudy (anonymous):

& everyone else too!